# LSCodec: Low-Bitrate and Speaker-Decoupled Discrete Speech Codec thanks:

### v1

@@ II-B Training Stage 1: Speech VAE
The decoder predicts two features: mel-spectrogram and SSL semantic tokens from the original un-perturbed content segment. In other words, the VAE model performs multi-task learning that simultaneously regresses towards acoustic mel features and classifies towards SSL semantic tokens in a fixed vocabulary size. The ground truth SSL tokens are extracted using HuBERT [9] and k-means clustering. The prediction of mel-spectrograms ensures reconstruction ability of the bottleneck features. As semantic tokens retain rich content information with a compact discrete space, the SSL token prediction task is crucial for guiding the bottleneck features to encode sufficient contents. Meanwhile, prosody information is largely damaged in the HuBERT tokens [12], so keeping the mel prediction task ensures prosody to be encoded in the bottleneck. Note that SSL models do not require labels to train, so the whole training pipeline is still unsupervised unlike FACodec [21].
@@ III-A Data and Model Setup
We use the output of the sixth layer from a pretrained WavLM-Large model [^4] https://github.com/microsoft/unilm/tree/master/wavlm to provide timbre features. In the SSL token prediction task, we use pretrained HuBERT-Large [^5] https://huggingface.co/facebook/hubert-large-ll60k and perform 2048-centroid k-means clustering on around a 83-hour subset of our training data. As this SSL feature resides in 50Hz, in the 25Hz version of LSCodec, we repeat each token by 2 times along the temporal axis before the Conformer decoder. The 80-bin cepstral-normalized mel-spectrograms for reconstruction also have a frame rate of 50Hz.
@@ III-B Speech Reconstruction
To evaluate the reconstruction ability of LSCodec against other low-bitrate speech tokens, we use the testset-B in [30] that contains 500 utterances from unseen speakers in LibriTTS test-clean split. Since low-bitrate speech coding cannot fully restore every detail in the original waveform, we emphasize on comparing the preservation of key information, especially content and prosody. We use word error rate (WER) to measure intelligibility and gross pitch error (GPE) to measure prosody preservation. WERs are computed using NeMo-ASR [^6] https://huggingface.co/nvidia/stt_en_fastConformer_transducer_large with the ground truth transcriptions. GPE stands for the ratio of pitch values larger than 20% relative error from the ground truth. We also conduct MOS tests as an subjective evaluation of audio quality. The baselines include semantic SSL tokens with CTX-vec2wavα vocoders, where wav2vec 2.0 [10] tokens are obtained from its inner quantizer, and HuBERT tokens are 2048-centroid cluster indexes from the last layer [^7] Same as that used in SSL token prediction task.. To enable a fair comparison among the low-bitrate acoustic speech codecs, we compare codecs with bitrate lower than 1kbps in Table I, i.e. EnCodec [5] and TiCodec [17] with only 1 VQ layer and WavTokenizer-small [25] in 40Hz. Official checkpoints are used for these baselines.
@@ IV Conclusion
We propose LSCodec, a low-bitrate and speaker-decoupled speech codec. LSCodec constructs a three-stage training framework with speaker perturbation. A VQ layer is applied after a VAE that disentangles speaker in a continuous space. A token vocoder is then trained upon the quantized codes. Experiments prove that LSCodec achieves strong reconstruction and VC ability given a bitrate as low as 0.25kpbs. Ablation studies validate the proposed training framework. LSCodec paves the way of future LM-based speech generation for its highly compact representation space. Explorations on stronger perturbation methods and better prosody preservation are valuable for future work.

### v2

@@ II-B Training Stage 1: Speech VAE
The decoder predicts two features: mel-spectrogram and SSL semantic tokens from the original un-perturbed content segment. In other words, the VAE model performs multi-task learning that simultaneously regresses towards acoustic mel features and classifies towards SSL semantic tokens in a fixed vocabulary size. The ground truth SSL tokens are extracted using WavLM [11] and k-means clustering. The prediction of mel-spectrograms ensures reconstruction ability of the bottleneck features. As semantic tokens retain rich content information with a compact discrete space, the SSL token prediction task is crucial for guiding the bottleneck features to encode sufficient contents. Meanwhile, prosody information is largely damaged in the discrete WavLM tokens [12], so keeping the mel prediction task ensures prosody to be encoded in the bottleneck. Note that SSL models do not require labels to train, so the whole training pipeline is still unsupervised unlike FACodec [21].
@@ III-A Data and Model Setup
We use the output of the sixth layer from a pretrained WavLM-Large model [^4] https://github.com/microsoft/unilm/tree/master/wavlm to provide timbre features. In the SSL token prediction task, we use the same pretrained WavLM and perform 2048-centroid k-means clustering on around a 83-hour subset of our training data. As this SSL feature resides in 50Hz, in the 25Hz version of LSCodec, we repeat each token by 2 times along the temporal axis before the Conformer decoder. The 80-bin cepstral-normalized mel-spectrograms for reconstruction also have a frame rate of 50Hz.
@@ III-B Speech Reconstruction
To evaluate the reconstruction ability of LSCodec against other low-bitrate speech tokens, we use the testset-B in [30] that contains 500 utterances from unseen speakers in LibriTTS test-clean split. Since low-bitrate speech coding cannot fully restore every detail in the original waveform, we emphasize on comparing the preservation of key information, especially content and prosody. We use word error rate (WER) to measure intelligibility and gross pitch error (GPE) to measure prosody preservation. WERs are computed using NeMo-ASR [^5] https://huggingface.co/nvidia/stt_en_fastConformer_transducer_large with the ground truth transcriptions. GPE stands for the ratio of pitch values larger than 20% relative error from the ground truth. We also conduct MOS tests as an subjective evaluation of audio quality. The baselines include semantic SSL tokens with CTX-vec2wavα vocoders, where wav2vec 2.0 [10] tokens are obtained from its inner quantizer, and WavLM tokens are 2048-centroid cluster indexes from the last layer [^6] Same as that used in SSL token prediction task.. To enable a fair comparison among the low-bitrate acoustic speech codecs, we compare codecs with bitrate lower than 1kbps in Table I, i.e. EnCodec [5] and TiCodec [17] with only 1 VQ layer and WavTokenizer-small [25] in 40Hz. Official checkpoints are used for these baselines.
@@ IV Conclusion
We propose LSCodec, a low-bitrate and speaker-decoupled speech codec. LSCodec constructs a three-stage training framework with speaker perturbation. A VQ layer is applied after a VAE that disentangles speaker in a continuous space. A token vocoder is then trained upon the quantized codes. Experiments prove that LSCodec achieves strong reconstruction and VC ability given a bitrate as low as 0.25kbps. Ablation studies validate the proposed training framework. LSCodec paves the way of future LM-based speech generation for its highly compact representation space. Explorations on stronger perturbation methods and better prosody preservation are valuable for future work.